# Shifting Attention to You: Personalized Brain-Inspired AI Models

### v1

@@ 3.5 Limitations and Need for Diverse Data
In thinking about the potential applications of personalized models, it is important to recognize the current limitations and the need for more diverse data. The success of our models in aligning with human cognition highlights the need for diverse and representative data. Incorporating behavioral and neural data from a broad participant pool reduces biases from homogeneous training sets, leading to more equitable and generalizable models (Born et al., 2024). However, beyond participant diversity, capturing the variability of real-world sensory experiences is equally important. As demonstrated in Figure 2B, CLIP-HBA-MEG, fine-tuned on group-level MEG data, exhibited strong alignment with human neural representations for clear, naturalistic stimuli but showed reduced generalizability on degraded visual datasets. This suggests that training exclusively on high-quality images may limit practicality and robustness. Since human perception naturally adapts to varied conditions—low lighting, occlusions, and noise—future datasets should reflect this variability by exposing the same participants to a range of visual conditions. This approach would enhance AI’s ecological validity, making models more adaptable and effective in real-world settings.
@@ 3.6 Conclusion
Our work underscores the versatility of the CLIP-HBA framework in adapting to a rich spectrum of data modalities, as demonstrated by its robust performance on large-scale, naturalistic stimulus collections. Looking ahead, systematic data collection efforts—ranging from active tasks within scanning environments to naturalistic interactions in immersive settings—hold the promise of revealing how human embeddings evolve when contextual demands change. By exploring active sensing, where representations directly support goal-directed actions, we may uncover entirely new dimensions of cognitive processing. The THINGS initiative (Hebart et al., 2019, 2023, 2020) has already paved the way for training models on diverse data types, and further endeavors that include neurodiverse populations and varied behavioral contexts will only enhance models like CLIP-HBA. Such efforts are not just a testament to the progress made so far but also a compelling challenge to further capture and understand the intricate dynamics of human perception and cognition.
@@ 4.1 Participants
Participants were drawn from existing datasets involving behavioral and neural measurements related to object recognition and mental representations. Behavioral data were obtained from a study that acquired extensive human similarity judgments of natural objects using the THINGS database. (Hebart et al., 2019, 2020). Neural data were sourced from magnetoencephalography (MEG) recordings collected in prior studies (Hebart et al., 2023) (Grootswagers et al., 2017a) (Grootswagers et al., 2017b) (Cichy et al., 2016). All participants provided informed consent in accordance with institutional guidelines approved by the relevant ethics committees.
@@ 4.2.1 THINGS Dataset and SPoSE Behavioral Embeddings
The THINGS dataset consists of a comprehensive compilation of 1,854 natural object concepts accompanied by high-quality images and semantic embeddings derived from the Sparse Positive Similarity Embedding (SPoSE) model (Hebart et al., 2019). The SPoSE model deduces 66 interpretable human behavioral embeddings based on similarity judgments, capturing the semantic dimensions that structure human mental representations of objects. We subsampled 1806 object stimuli with their corresponding SPoSE embedding for the training and testing of CLIP-HBA-Behavior, excluding 48 fully-sampled stimuli data for measuring the model’s behavioral alignment.
@@ 4.2.2 MEG Data
We utilized two separate MEG datasets to fine-tune CLIP-HBA-MEG at both group and individual levels (Table 2). For group-level training, we subsampled averaged MEG RDMs from three participants and selected 1,806 out of 1,854 visual stimuli from the THINGS MEG dataset (Hebart et al., 2023), reserving the same 48 fully-sampled stimuli for behavioral validation. This dataset was chosen for its relatively large stimulus set, ensuring generalizability and providing a robust foundation for demonstrating the model’s ability to learn meaningful neural representations. For individual-level training, we used a dataset with 118 stimuli but a larger sample of 15 participants (Cichy et al., 2016), prioritizing the number of participants to assess the model’s capacity for individualized neural tuning and capturing inter-individual differences. To maintain consistency across training conditions and computational efficiency, we downsampled the individual-level dataset to 200Hz, matching the sampling rate used in group-level training.
@@ 4.3.1 Baseline CLIP Model
We employed the Contrastive Language–Image Pre-training (CLIP) model as our base architecture (Radford et al., 2021). CLIP integrates a visual encoder, built on the Vision Transformer (ViT) architecture (Dosovitskiy et al., 2020), with a text transformer encoder. Pre-trained on a large dataset of image-text pairs, the model aligns visual and textual representations within a shared embedding space. For this study, we specifically employed CLIP-ViT-L/14, a variant with an increased parameter count to enable richer representational learning. This same model configuration serves as our baseline, ensuring a fair comparison when evaluating performance enhancements.
@@ 4.3.2 Fine-Tuning with Behavioral Data
To balance fine-tuning effectiveness with computational efficiency, we employed Weight-Decomposed Low-Rank Adaptation (DoRA) (Liu et al., 2024) modules into the out-projection of the attention layers, specifically in the final text encoder layer and the last two vision encoder layers. DoRA extends the widely used parameter-efficient fine-tuning (PEFT) method, LoRA (Hu et al., 2021), by decomposing pre-trained weights into separate magnitude and directional matrices. This decomposition enables distinct magnitudinal and directional learning during fine-tuning, closely replicating the optimization dynamics of training directly on the original model weights. Only the DoRA parameters are updated, while the original layers remain frozen, significantly reducing the number of necessary training parameters and computational demands. Furthermore, to leverage CLIP’s pre-training on large image-text datasets, we limit the fine-tuning updates to only the last few layers of text and vision encoders, preserving transfer learning capabilities while maintaining the pre-trained model’s inherent generalizability (Fahes et al., 2024).
@@ 4.3.3 Behavioral Predictions Evaluation
We evaluated the fine-tuned CLIP-HBA-Behavior model’s ability to predict human behavior using a triplet odd-one-out task on a held-out portion of the data of 48 sample objects. (Hebart et al., 2020). This set was fully sampled to cover a wide range of naturalistic objects, and these 48 object images were specifically excluded during training to prevent any data leakage. Representational dissimilarity matrices (RDMs) were constructed from the model’s embeddings by calculating pairwise cosine distances (Kriegeskorte, 2008). These RDMs were compared to RDMs derived from actual behavioral choices using Spearman rank-order correlations (Zar, 1972), assessing the degree of alignment between model predictions and human judgments.
@@ 4.3.4 Dynamic Neural Fine-Tuning
Stage 1 - Feature Reweighting: In this phase, the ViT parameters remain frozen while only the feature reweighting matrix $\mathbf{W}$ is optimized. To enforce temporal consistency, dependencies between neighboring time points are introduced through an average sliding window. Optimization is performed using the AdamW optimizer (Loshchilov and Hutter, 2019) for the initial epochs until convergence. This step establishes foundational temporal layer dependencies without altering the underlying visual processing capabilities of the pre-trained CLIP model.
@@
The model incorporates temporal modulation through several key components: 1) visual feature magnitude scaling ($\alpha_{T}$) controlled by neural response richness (Carlson et al., 2011) (Carlson et al., 2013), and 2) semantic binding strength ($\beta_{T}$) determined by time generalization patterns (Tovar et al., 2020b) (King and Dehaene, 2014). Additionally, to better mimic the biological nature of human perception—where attention is never fully fixed on a stimulus while ignoring all else—we introduce a mechanism that allows the model to "think of something else" by incorporating dimension-specific Gaussian noise during semantic binding. This ensures a degree of randomness, resembling the natural variability in human neural processing:
@@
MEG decoding RDMs serve as the target representations within this pipeline. Since neural data lacks a directly comparable 66-dimensional embedding space like the behavioral SPoSE embedding, we transform the model’s output embeddings into representational dissimilarity matrices (RDMs) using Pearson correlation (Kirch, 2008). This conversion translates the model embeddings into a common representational space with neural data, allowing alignment comparison against the target MEG decoding RDMs.
@@ 4.3.5 Neural Data Alignment Evaluation
We validated the efficacy of our neurally fine-tuned CLIP-HBA-MEG model through both behavioral and neural alignment analyses. For the behavioral alignment evaluation, we used spearman correlation to compare every slice of the model’s dynamic RDMs against the static behavioral RDM derived the same 48 sampled stimuli, excluded from the neural fine-tuning process. For the neural alignment evaluation, we assessed the model’s performance by comparing its timepoint-to-timepoint RDMs with average-participant MEG decoding RDMs from the THINGS MEG dataset and all mentioned additional validation datasets mentioned in Table 3, using Spearman rank correlation (Zar, 1972).
@@ 4.5.1 Dynamic Saliency Mapping of CLIP-HBA-MEG
We achieved millisecond-level visualization of dynamically important regions within any given visual stimulus using the fine-tuned CLIP-HBA-MEG model and Randomized Input Sampling for Explanation (RISE) (Petsiuk et al., 2018). The original RISE method identifies critical regions for model processing by comparing classification accuracy between unmasked images and randomly masked inputs.

### v2

@@ 3.5 Limitations and Need for Diverse Data
In thinking about the potential applications of personalized models, it is important to recognize the current limitations and the need for more diverse data. The success of our models in aligning with human cognition highlights the need for diverse and representative data. Incorporating behavioral and neural data from a broad participant pool reduces biases from homogeneous training sets, leading to more equitable and generalizable models [9]. However, beyond participant diversity, capturing the variability of real-world sensory experiences is equally important. As demonstrated in Figure 2B, CLIP-HBA-MEG, fine-tuned on group-level MEG data, exhibited strong alignment with human neural representations for clear, naturalistic stimuli but showed reduced generalizability on degraded visual datasets. This suggests that training exclusively on high-quality images may limit practicality and robustness. Since human perception naturally adapts to varied conditions—low lighting, occlusions, and noise—future datasets should reflect this variability by exposing the same participants to a range of visual conditions. This approach would enhance AI’s ecological validity, making models more adaptable and effective in real-world settings.
@@ 3.6 Conclusion
Our work underscores the versatility of the CLIP-HBA framework in adapting to a rich spectrum of data modalities, as demonstrated by its robust performance on large-scale, naturalistic stimulus collections. Looking ahead, systematic data collection efforts—ranging from active tasks within scanning environments to naturalistic interactions in immersive settings—hold the promise of revealing how human embeddings evolve when contextual demands change. By exploring active sensing, where representations directly support goal-directed actions, we may uncover entirely new dimensions of cognitive processing. The THINGS initiative [18, 19, 25] has already paved the way for training models on diverse data types, and further endeavors that include neurodiverse populations and varied behavioral contexts will only enhance models like CLIP-HBA. Such efforts are not just a testament to the progress made so far but also a compelling challenge to further capture and understand the intricate dynamics of human perception and cognition.
@@ 4.1 Participants
Participants were drawn from existing datasets involving behavioral and neural measurements related to object recognition and mental representations. Behavioral data were obtained from a study that acquired extensive human similarity judgments of natural objects using the THINGS database. [18, 25]. Neural data were sourced from magnetoencephalography (MEG) recordings collected in prior studies [19] [27] [28] [11]. All participants provided informed consent in accordance with institutional guidelines approved by the relevant ethics committees.
@@ 4.2.1 THINGS Dataset and SPoSE Behavioral Embeddings
The THINGS dataset consists of a comprehensive compilation of 1,854 natural object concepts accompanied by high-quality images and semantic embeddings derived from the Sparse Positive Similarity Embedding (SPoSE) model [18]. The SPoSE model deduces 66 interpretable human behavioral embeddings based on similarity judgments, capturing the semantic dimensions that structure human mental representations of objects. We subsampled 1806 object stimuli with their corresponding SPoSE embedding for the training and testing of CLIP-HBA-Behavior, excluding 48 fully-sampled stimuli data for measuring the model’s behavioral alignment.
@@ 4.2.2 MEG Data
We utilized two separate MEG datasets to fine-tune CLIP-HBA-MEG at both group and individual levels (Table 2). For group-level training, we subsampled averaged MEG RDMs from three participants and selected 1,806 out of 1,854 visual stimuli from the THINGS MEG dataset [19], reserving the same 48 fully-sampled stimuli for behavioral validation. This dataset was chosen for its relatively large stimulus set, ensuring generalizability and providing a robust foundation for demonstrating the model’s ability to learn meaningful neural representations. For individual-level training, we used a dataset with 118 stimuli but a larger sample of 15 participants [11], prioritizing the number of participants to assess the model’s capacity for individualized neural tuning and capturing inter-individual differences. To maintain consistency across training conditions and computational efficiency, we downsampled the individual-level dataset to 200Hz, matching the sampling rate used in group-level training.
@@ 4.3.1 Baseline CLIP Model
We employed the Contrastive Language–Image Pre-training (CLIP) model as our base architecture [17]. CLIP integrates a visual encoder, built on the Vision Transformer (ViT) architecture [50], with a text transformer encoder. Pre-trained on a large dataset of image-text pairs, the model aligns visual and textual representations within a shared embedding space. For this study, we specifically employed CLIP-ViT-L/14, a variant with an increased parameter count to enable richer representational learning. This same model configuration serves as our baseline, ensuring a fair comparison when evaluating performance enhancements.
@@ 4.3.2 Fine-Tuning with Behavioral Data
To balance fine-tuning effectiveness with computational efficiency, we employed Weight-Decomposed Low-Rank Adaptation (DoRA) [51] modules into the out-projection of the attention layers, specifically in the final text encoder layer and the last two vision encoder layers. DoRA extends the widely used parameter-efficient fine-tuning (PEFT) method, LoRA [52], by decomposing pre-trained weights into separate magnitude and directional matrices. This decomposition enables distinct magnitudinal and directional learning during fine-tuning, closely replicating the optimization dynamics of training directly on the original model weights. Only the DoRA parameters are updated, while the original layers remain frozen, significantly reducing the number of necessary training parameters and computational demands. Furthermore, to leverage CLIP’s pre-training on large image-text datasets, we limit the fine-tuning updates to only the last few layers of text and vision encoders, preserving transfer learning capabilities while maintaining the pre-trained model’s inherent generalizability [53].
@@ 4.3.3 Behavioral Predictions Evaluation
We evaluated the fine-tuned CLIP-HBA-Behavior model’s ability to predict human behavior using a triplet odd-one-out task on a held-out portion of the data of 48 sample objects. [25]. This set was fully sampled to cover a wide range of naturalistic objects, and these 48 object images were specifically excluded during training to prevent any data leakage. Representational dissimilarity matrices (RDMs) were constructed from the model’s embeddings by calculating pairwise cosine distances [26]. These RDMs were compared to RDMs derived from actual behavioral choices using Spearman rank-order correlations [54], assessing the degree of alignment between model predictions and human judgments.
@@ 4.3.4 Dynamic Neural Fine-Tuning
Stage 1 - Feature Reweighting: In this phase, the ViT parameters remain frozen while only the feature reweighting matrix $\mathbf{W}$ is optimized. To enforce temporal consistency, dependencies between neighboring time points are introduced through an average sliding window. Optimization is performed using the AdamW optimizer [55] for the initial epochs until convergence. This step establishes foundational temporal layer dependencies without altering the underlying visual processing capabilities of the pre-trained CLIP model.
@@
The model incorporates temporal modulation through several key components: 1) visual feature magnitude scaling ($\alpha_{T}$) controlled by neural response richness [56] [57], and 2) semantic binding strength ($\beta_{T}$) determined by time generalization patterns [58] [59]. Additionally, to better mimic the biological nature of human perception—where attention is never fully fixed on a stimulus while ignoring all else—we introduce a mechanism that allows the model to "think of something else" by incorporating dimension-specific Gaussian noise during semantic binding. This ensures a degree of randomness, resembling the natural variability in human neural processing:
@@
MEG decoding RDMs serve as the target representations within this pipeline. Since neural data lacks a directly comparable 66-dimensional embedding space like the behavioral SPoSE embedding, we transform the model’s output embeddings into representational dissimilarity matrices (RDMs) using Pearson correlation [60]. This conversion translates the model embeddings into a common representational space with neural data, allowing alignment comparison against the target MEG decoding RDMs.
@@ 4.3.5 Neural Data Alignment Evaluation
We validated the efficacy of our neurally fine-tuned CLIP-HBA-MEG model through both behavioral and neural alignment analyses. For the behavioral alignment evaluation, we used spearman correlation to compare every slice of the model’s dynamic RDMs against the static behavioral RDM derived the same 48 sampled stimuli, excluded from the neural fine-tuning process. For the neural alignment evaluation, we assessed the model’s performance by comparing its timepoint-to-timepoint RDMs with average-participant MEG decoding RDMs from the THINGS MEG dataset and all mentioned additional validation datasets mentioned in Table 3, using Spearman rank correlation [54].
@@ 4.5.1 Dynamic Saliency Mapping of CLIP-HBA-MEG
We achieved millisecond-level visualization of dynamically important regions within any given visual stimulus using the fine-tuned CLIP-HBA-MEG model and Randomized Input Sampling for Explanation (RISE) [31]. The original RISE method identifies critical regions for model processing by comparing classification accuracy between unmasked images and randomly masked inputs.